the width of the rectangle is one-fourth of its length. if the perimeter is 40cm, find the dimensions of the rectangle.
Answer by ptaylor(2198)

Perimeter=2*Width+2*Length or P=2L+2W
Let L=the length of the rectangle
Then (1/4)L=width of the rectangle
So, our equation to solve is:
40=2*L+2*(1/4)L=2L+(1/2)L=2.5L divide each side by 2.5
L=16cm-----------------------Length of rectangle
W=(1/4)*16=4cm----------------Width of rectangle
CK
2*16+2*4=40
32+8=40
40=40
Hope this helps----ptaylor